OpenStudy (purple_pink):

Alan estimated that he would score 140 points the next time he bowled. He actually scored 152 points. What was the percent error in Alan’s estimate? Round to the nearest tenth of a percent, if needed. A. 6% B. 7.9% C. 8.6% D. 12%

OpenStudy (kittiwitti1):

I'm not exactly sure if this is the right way to do it, but I'd • subtract 140 from 152, which gets you the points he scored OVER his predicted score (12). • make a fractional ratio out of the value from previous step and 140 (12/140). simplify it until you get a decimal • convert that decimal to percentage form\[152-140=12\rightarrow\frac{12}{140}\rightarrow\frac{3}{35}\approx 0.086\text{ or }8.6\text{%}\]

OpenStudy (purple_pink):

woah! thanks! :D

OpenStudy (kittiwitti1):

I don't know if this is the correct method, though. I'd ask another person just to make sure ♦

OpenStudy (purple_pink):

Sure

OpenStudy (kittiwitti1):

Also, random trivia: if you look at the last four letters in your question's URL ... ☻

OpenStudy (purple_pink):

yes?

OpenStudy (kittiwitti1):

Just look at it: http://openstudy.com/study#/updates/56e86a50e4b093c37491dead

OpenStudy (kittiwitti1):

lol ☻

OpenStudy (purple_pink):

oh loooooool

OpenStudy (kittiwitti1):

Well, I asked a math pro on here to check my work. Looks like I did it right. Thanks @agent0smith ♥ Good luck on your studies, @purple_pink ♣

OpenStudy (purple_pink):

thank you sooo much do you mind helping me with 2 more?

OpenStudy (kittiwitti1):

Sure, just post a new question ♪

OpenStudy (purple_pink):

i know thanks!

OpenStudy (kittiwitti1):

No problem ♫